OpenStudy (anonymous):

How is it possible to simplify: cos(alpha)(-1) possibly using an identity?

OpenStudy (perl):

it is hard to say what this represents , is this an equation or inverse cosine

OpenStudy (anonymous):

this is the last part of my question all I need is to simplify unless this is the most it may be simplified

OpenStudy (anonymous):

\[\cos(180+\alpha)\]

OpenStudy (perl):

we can simplify that,

OpenStudy (perl):

you can actually apply the cos angle sum formula cos (x + y) = cos(x) * cos(y) - sin(x)*sin(y)

OpenStudy (anonymous):

how.... I don't want a final answer, just help on getting there.

OpenStudy (perl):

cos(180 + a) = cos(180)*cos(a) - sin(180)*sin(a)

OpenStudy (anonymous):

okay, and I have. It's suppose to be done in terms of alpha

OpenStudy (perl):

a = alpha

OpenStudy (anonymous):

i understood what you ment I just didn't want to go to the specific menu....

OpenStudy (perl):

yeah i hate the latex menu myself :D

OpenStudy (anonymous):

the question states to answer it in terms of Alpha and i came to: \[\cos(\alpha)(-1)\]

OpenStudy (perl):

right, do you see why?

OpenStudy (anonymous):

yes, but is it possible to simplify or is that the final step?

OpenStudy (perl):

cos(180 + a) = cos(180)*cos(a) - sin(180)*sin(a) = (-1)*cos(a) - 0 * sin(a) = -cos(a)

OpenStudy (anonymous):

ohhhhh okay. I understand :) thanks